OpenStudy (anonymous):

Perform any indicated operations and simplify (extract as much from roots as possible): (sqrt(24x^3y^2))(sqrt(3x^3y)) Please show step by step

OpenStudy (anonymous):

\[\sqrt{24x ^{3}y ^{2}}\sqrt{3x ^{3}y}\]

OpenStudy (marco26):

\[\sqrt{4*2*3x^2xy^2} \sqrt{3x^2xy}=2xy \sqrt{6x}[(x)\sqrt{3xy}]=2x^2y \sqrt{18x^2y}=2x^2y \sqrt{9*2x^2y}=6x^3y \sqrt{2y}\]